### v1

# Artificially intelligent agents in the social and behavioral sciences

Abstract
We review the historical development and current trends of artificially intelligent agents (agentic AI) in the social and behavioral sciences—from the first programmable computers, and social simulations soon thereafter, to today’s experiments with large language models. This overview emphasizes the role of AI in the scientific process and the changes brought about, both through technological advancements and the broader evolution of science from around 1950 to the present. Some of the specific points we cover include: the challenges of presenting the first social simulation studies to a world unaware of computers, the rise of social systems science, intelligent game theoretic agents, the age of big data and the epistemic upheaval in its wake, and the current enthusiasm around applications of generative AI, and many other topics. A pervasive theme is how deeply entwined we are with the technologies we use to understand ourselves.
1 Introduction
For millennia, humanity has cherished a dream of creating machines that can think and behave like ourselves [124]. The closer we get to realizing that goal, the more familiar the image in the mirror appears, and the stronger we believe that understanding the image is really understanding ourselves. The idea of using artificial human replicas in the social and behavioral sciences has seemed self-evident to scientists throughout the ages—every major related technological breakthrough has soon been followed by applications to research on humans. At the time of writing, advances in artificial intelligence, with buzzwords like generative AI (GenAI) and large language models (LLMs), are frequent headline news even outside of science. No wonder this has created yet a wave of research aiming to understand social and behavioral phenomena through new technologies [195, 67, 12, 229]. However, this time might be different. For the first time, we have access to “strong AI” agents [185] that seemingly can take almost any of our roles in experiments and observations (Fig. 1). Then again, this new technology changes our society and behavior in a feedback loop that obscures the road ahead. This paper attempts to cover the developments over the last seventy-five years, leading up to the present use of AI as a research tool. We will emphasise that there has been a bidirectional flow of ideas between AI and the social/behavioral sciences ever since the inception of the former, and try to rectify the picture of AI as a tool arriving to the social/behavioral sciences from afar [55].
Telling the story of AI and the social/behavioral sciences is challenging not only because of the entwined flows of ideas and influence, but also because of the many roles AI can have in it. First, as alluded to above, one could see AI as a building block of the scientific method, as an analysis tool, or as a digital twin technology (i.e., black box rendition of people or institutions), etc. Second, one could approach AI as an object of study and try to understand, for example, how behaviorism, which influenced Turing’s test [208, 24], transpires in LLMs and generative AI. Third, one could see the emergence of AI as a social phenomenon, where the imagination of investors, futurists, and scientists has aligned to bring us to where we are today. Finally, one could view AI as a social force changing both the interaction between people and the people themselves [224]. The availability of new research tools aside, to understand this emergent human-machine society is the most pressing motivation for experimenting with agentic AI at the time of writing [166, 158, 206, 207]. In Fig. 1, we sketch some approaches in the literature.
Figure 1: Illustation of knowledge-building information flows contributing to the behavioral and social sciences involving humans and AI. Panel A illustrates how AI has always been modeled on human intelligence. This has also shaped our understanding of ourselves. B shows the typical simulation setup of AI agents mimicking a social situation. In C, AI aids in collecting data. Panel D shows experiments with AI agents interacting with humans. Panel E illustrates an AI scientist building knowledge without human intervention. Even though AI-generated science exists today, this particular scenario might yet be fictitious, especially considering the ethical aspects of AI probing human life through interventions [150].
The main thesis of this paper is that AI’s contribution to the social and behavioral sciences has been highly complex, driven by a multitude of processes, two of which stand out. First, any technical breakthrough is typically rapidly adopted by open-minded social and behavioral scientists, who use it to design new types of investigations. Second, these lines of research branch off to follow a slower-paced evolution centered around the scientific content. For this reason, it is hard to present the historical developments as a timeline. Rather, there are several computational or AI-infused streams running in parallel. We cover these roughly by their starting date, and note that they are all still active research areas to this date.
Before proceeding to describe the history and present of AI in the social and behavioral sciences, let us say a few words about what this paper is and is not. First, not all the historical approaches we present are based on intelligent agents; however, they all help to set the direction, scope, expectations, and methodologies of agentic-AI science as it stands today. Second, we will not give a comprehensive review of the recent literature as other papers do that [55], but we will sketch the current state and trends. Finally, even though there is an overlap between the behavioral and cognitive sciences, we will not cover that interface very deeply—that is indeed a rich and fascinating topic in its own right [138, 136, 137, 24], but more philosophical compared to our paper.
2 Simulations and early agent-based models
It might seem surprising at first, just how soon after the arrival of the first programmable computers, they were brought into use to simulate humans. An explanation is that the scientists behind these applications were, to a great extent, a part of the same intellectual movement that led to the development of computers themselves. By the time electronic computers became accessible for academic research, in the first few years of the 1950s, key ideas of simulational research were already public. For example, the idea that people’s roles can be characterized by how they are connected to others [140] and the thought that social agents could best be modeled as part random, part purpose-driven were established [167]. Furthermore, how to analyze the output from computer simulations was, of course, known from experimental and observational science. So even if the mode of scientific explanation manifested by computer simulations—famously described by James Coleman’s boat diagram [35, 225]—has taken a long time to become accepted by the mainstream, the elements were already there when the first computers became available to science. The transition was indeed so smooth that there were papers based on stochastic simulations with manual random number generation, like Helen Abbey’s study of the Reed-Frost epidemic model from 1952 [1]. [^1] For a more celebrated example, jumping ahead to 1969, stochastic modeling without the use of computers was also the premise for Thomas Schelling’s segregation model [180]. In the wake of the civil rights movement and the heightened racial tensions of the era, Schelling used a rule-based model of individuals moving their domiciles to make the point that people don’t have to have a strong racial bias for segregation to happen—it can emerge from the migration itself.
Early computer-aided studies in the behavioral and social sciences, of course, faced some problems that we do not have today—this was a time when computers were something people might have read about in a newspaper but usually never even seen in their everyday lives. In 1966, authors could feel compelled to explain computers as [61]:
…a clerk with several virtues but one which is no better than its instructions. To appreciate the way in which this clerk operates, one should realize that a computer is, in principle, nothing more than a desk calculator and a note pad.
With describing these transistorized clerks out of the way, the next challenge was to convince the readers of their methodological merits. In this aspect, the papers have a much more modern feel. We can notice five such themes:
Nothing new
Several early papers point out that simulating humans did not start with digital computers. Harry Harman wrote in a military white paper in 1961 that [73]:
Simulation may be traced back to the beginning of time—be it the make-believe world of the child at play, or the adult make-believe world of the stage.
In his The Sciences of the Artificial, Herbert Simon makes the same point by referring to an analog simulator of Keynesian economics (the Moniac[192].
To facilitate systems thinking
As we will discuss further below in the context of systems science, one radical insight in 20th-century science was that when a system contains feedback loops, so things affect themselves, one cannot discuss causality in usual terms [101]. Mathematical models get hard to solve, but, as some early papers point out, could still be analyzed by computer simulations [68, 226].
To do experiments otherwise impossible
This was perhaps the most common motivation and one that has fallen somewhat out of fashion. Quoting the psychologist Kenneth Colby (1967) [34]:
Before the computer program we had no satisfactory approach to huge, complex, ill-defined systems difficult to grapple with, not only because of their multivariate size but also because of a property of elusiveness which in psychology is mainly a function of vagueness in that the limits of inclusiveness of conventional terms are unclear.
The “complex, ill-defined system” Colby had in mind was nothing less than a “neurotic patient.” For another example, de Sola Pool and Kessler use this motivation in their essay “The Kaiser, the Tsar and the Computer” to present their simulation-based decision support system for international political conflicts, Crisiscom [40].
Proof-of-concept models
Another motivation for computer simulations in the 1950s and ’60s was that they could help validate qualitative theorizing. What good is a theory if it cannot even reproduce qualitative results? AI pioneer Edward Feigenbaum wrote in his 1963 essay “Computer simulations of human behavior” [48]:
One of the advantages of computer simulation is this one, of guaranteeing sufficiency and completeness. The computer simulation model will not operate if you forget anything. If you fail to take into account some necessary mechanism that, in the verbal description of a theory, you might readily pass over, the computer simulation will not run. Thus, one is forced to focus attention on all of the mechanisms—those that are not well understood as well as those that are well understood—and all of the subtle effects and interactions which are implicit in one’s model. In a sense, this business of accounting for all of the information processing is a very powerful mental discipline.
For forecasting and scenario testing
Forecasting and scenario testing are, of course, areas where simulations could inform decision-makers. The above-mentioned crisis simulator by de Sola Pool is a good example and may be the reason political science was very early to jump on the simulation bandwagon [40]. Another example by the same authors was a simulation program for American presidential elections [39], which, at the time, played a similar role to that of election predictors and prediction markets today [222].
3 Systems and information
Another side of the intellectual and scientific trends that led to the development of the electronic computer was a focus on systems—collections of units that cannot be understood by reducing them to simple causal relations [101, 79, 8, 220]. One academic movement based on this tenet, predating artificial intelligence, was cybernetics. Cybernetics encompassed topics such as control, communication, information transfer, and an aspiration to not only transcend scientific disciplines but also break the barriers between the applied and theoretical sciences.
Cybernetics influenced the social and behavioral sciences in several waves. The first wave was centered around ten so-called Macy conferences [160] in 1946–53. Some of these meetings had an explicit social theme. For example, the second conference gathered sociologists Paul Lazarsfeld, Talcott Parsons, and Robert Merton, psychologist Kurt Lewin, and anthropologists Margaret Mead and Gregory Bateson, to discuss teleological mechanisms in society. Here, “teleological” referred to Rosenblueth, Wiener, and Bigelow’s then-recent essay “Behavior, purpose and teleology” [171], which argued for cybernetics as an alternative to behaviorism. [^2] The cyberneticians saw behaviorism as overly simplistic in its inablity to handle causal feedback loops. Even if the first generation of cyberneticians didn’t directly inspire computational approaches to the social sciences, their systems thinking had an impact on, e.g., Parsons’ [156] and Merton’s [133] work from around this time.
It is hard to pinpoint the first systems-oriented computational framework, but the most successful of the early ones was arguably Jay Forrester’s systems dynamics from the late 1950s [51]. Inspired by the management of military operations, Forrester proposed coupled differential equations with delays as a universal systems-simulation paradigm and went on to make an unreplicated political impact. In the early 1970s, the environmentalist movement had its first peak, the public belief in computer technology was strong, and systems thinking was making an academic impact [149]. All this set the stage for the report Limits to Growth [129] that presented systems dynamics simulations of the global population, food supply, and environmental pollution to argue that the economic growth paradigm is doomed to failure in our world of limited resources [182]. Limits to Growth had a tremendous political impact, and one can argue that today’s degrowth movement [94], the UN’s sustainable development goals, and academic research on Earth system analysis can all be traced back to it [182].
Figure 2: An illustration of trace-data to support governance adapted from Stafford Beer’s 1975 A Platform for Change [17]. Even though Beer envisioned a planning economy capable of catering to the needs of the individual via data mining and cybernetic control [130], this figure inadvertently seems to anticipate today’s information economy and digital surveillance [116, 95].
In the 1960s, cybernetics had a computerized revival, especially in Britain [162]. Notably, Stafford Beer and colleagues presented cybernetic support tools for governance that foreboded today’s information economy and computational social science, see Fig. 2. Beer also almost got the chance to put his theories to work in the socialist regime of Chile’s Salvador Allende, where he led the development of the governance support system Cybersyn [130]. [^3] Allende’s regime came to a dramatic end with the 1973 coup d’état, but interestingly, the new regime under General Augusto Pinochet also played a minor part in our story. The military leadership consulted the Nobel laureate and proponent of laissez-faire economics, F. A. Hayek, to rebuild the Chilean economy—the same Hayek who helped lay the foundations for the currently prevailing connectionist approach to AI (see below, and Ref. [74]).
Figure 3: Part of the flow chart of Newell and Simon’s General Problem Solver [144]—the part related to “means-end analysis.” Without going into detail about how the constituents are manifested in the program, we can still appreciate how different Newell and Simon’s approach is from today’s machine-learning-based AI. Newell and Simon explicitly stated that Turing was not an inspiration for their AI [125], which seems evident from the General Problem Solver.
4 Pioneers of artificial intelligence
Even though many foundational works were published earlier—like Turing’s “Computing machinery and intelligence” [208] and the neuronal model of McCulloch and Pitts [126]—the academic field of artificial intelligence usually sets its birth date to 1956. This was the year when John McCarthy and Marvin Minsky organized a research program titled “Dartmouth Summer Research Project on Artificial Intelligence.” The term artificial intelligence was allegedly a political choice to stand out from neighboring fields [125]. [^4] Indeed, several of the fields and subfields mentioned in this article are primarily social rather than topical constructs.
Throughout its history, AI has included a wide variety of approaches to machine intelligence. It has also had many ups and downs, and the term’s connotations have fluctuated in concert with the paradigm of the day. Already from the onset, some researchers, like Allen Newell and Herbert Simon, approached AI by reverse engineering higher-order cognitive abilities like reasoning and problem-solving; others, like Walter McCulloch and Frank Rosenblatt, focused on the computational capabilities of connected nerve cells. These roughly correspond to the symbolic and connectionist approaches to AI [139, 174].
4.1 Simon et al.’s symbolic reasoning AI
Herbert Simon’s career started with the study of administrative behavior and organizational decision-making [190]. Although Simon’s output was extremely diverse, one can often trace his ideas back to organizational decision processes of one sort or another [193]. At the aforementioned Dartmouth conference, Simon and his long-standing collaborator Allen Newell presented their theorem-proving program “The Logic Theorist” that was capable of solving logical propositions and is sometimes regarded as the first AI program [125]. Continuing this line, Newell and Simon built the General Problem Solver (GPS)—a program simulating human reasoning in general terms [144] (see Fig. 3).
Simon was later awarded the Nobel Economics Prize primarily for his work on bounded rationality [191]. However, Simon’s social science most pertinent to this paper was presented in The Sciences of the Artificial [192], which was foundational to both design theory and the study of complex social systems. One often misunderstood aspect of this work today is that Simon called for an investigation of the design process of artificial objects and their de facto usage, not just their functionality [192]:
The thesis is that certain phenomena are “artificial” in a very specific sense: they are as they are only because of a system’s being molded, by goals or purposes, to the environment in which it lives. If natural phenomena have an air of “necessity” about them in their subservience to natural law, artificial phenomena have an air of “contingency” in their malleability by environment.
While the AI of Simon et al. may seem a bit outdated today, it did have an immense impact, much in the form of “expert systems” [43]. Expert systems constitute perhaps the first commercially successful type of AI application, with users primarily in corporate and institutional management. Their advances in knowledge representation linger to this day in fields like knowledge management and organizational learning [146].
In a prescient work [30], Kathleen Carley and colleagues attempted to connect Allen Newell’s “Soar” architecture for a general-purpose AI (or cognitive architecture as they termed it) [143] into a model of organizational decision making. Carley et al. considered agents who need to find components in a warehouse to complete incoming orders, and require both coordination and communication to achieve this. See Table 1 for a list of features needed to replicate a human in a social simulation according to Ref. [30] (a topic Carley and Newell expanded on in Ref. [31]).
Table 1: Required capabilities of social agents according to Carley and Newell [30]. The authors admit that they have managed to achieve all except the ones pertaining to “social analysis.”
Perception and Action
      Perceives the environment
      Physically manipulates objects
      Moves self to different locations
Memory
      Location
      People
      Task
Instruction
      Can be incomplete
Task Analysis
      Decomposes task
      Coordinates subtasks for self to do
Communication Skills
      Asks questions / Provides answers
      Gives commands / Receives commands
      Talks to a single individual / Talks to a group
Social Analysis
      Models of other agents
      Model of organization
4.2 The early connectionists
The other major branch of early artificial intelligence, connectionism, was the antecedent of many of the most hyped-up technologies of today [174]. This is the tradition of building machine intelligence on networks of model neurons. The starting point of connectionism was the perceptron—a neuronal model invented by McCulloch and Pitts in 1943 [126]. Fifteen years later, Frank Rosenblatt implemented a network of perceptrons with custom-made hardware [170]. Rosenblatt’s paper is an interesting read, going far beyond technicalities and crediting, among others, the above-mentioned F. A. Hayek for his philosophically trailblazing The Sensory Order [74]. After their initial success, artificial neural networks entered a hiatus that would last throughout the 1970s [49].
5 Embodied cognition and autopoiesis
The approaches to artificial intelligence and modeling of humans described above (except, perhaps, the perceptron) have in common the fact that they do not depend on a physical body. They are consistent with the Cartesian dictum that the mind and the body can be considered separate, with the mind operating solely as an information-processing device. This orthodoxy has been challenged several times: Merleau-Ponty’s phenomenology of perception in the 1940s [132], Gibson’s ecological psychology of the 1960s [59], and by the early 1990s, embodied cognition [209, 221]. The latter also became influential in robotics, where, logically, artificial intelligence should be localized to the robot itself, in its environment [54, 159].
From the same spirit of skepticism as above—maybe things that were considered separate actually are sufficiently connected to be considered as one—cybernetics had another energy injection in the early 1970s with second-order cybernetics making the point that an observer cannot be fully detached from the observed object or system [128]. It is not a coincidence that this line of thinking coincided with the ever more inclusive system simulation models (e.g., The Limits to Growth[149, 129, 169]. Furthermore, the inevitable consequence of allowing wider and looser boundaries is that one’s study objects will have more capabilities. Against this backdrop, Humberto Maturana and Francisco Varela arrived at their concept of autopoiesis in 1972 [122]. This refers to the ability of living systems to reproduce themselves, by organizing, assembling, and, to some extent, producing their constituents. Autopoiesis has had a broad influence—from manufacturing engineering, via the philosophy of mind, to sociology. In the latter, thinkers like Niklas Luhmann [115] and Edgar Morin [141] explored the autopoietic nature of social systems.
Around the turn of the millennium, fields like robotics and human-computer interaction expanded into topics like the integration of socially aware agents in human populations (see, e.g., Ref. [38]). This stream of research was strongly influenced by embodied cognition. Furthermore, it approximated today’s discussion of AI in society, and even made some points that have since been lost. For example, this research stressed the importance of embodied perspectives and narrative understandings of courses of events if one seeks to make AI capable of replacing humans. This literature saw designing a robot friend beyond a simulacrum [23] as a distant goal, which, at the time of writing, is no longer on the horizon.
Figure 4: An illustration of the Sugarscape model by Epstein and Axtell [44]—an archetypal agent-based model. The simulation is based on agents spread out over a grid and an unevenly distributed natural resource—the “sugar.” A darker shade of the background color indicates a higher sugar concentration. In principle, one could use any grid geometry and sugar distribution, but Sugarscape has become synonymous with the shown bimodal distribution on a 50 $\times$ 50 grid. Agents move to optimize a utility function within their “field of vision”—the gray surroundings of the white dots in panel A. If direct sugar consumption is all the agents care about, their population distribution will come to resemble the sugar distribution (panel A). By adding abilities to the agents and terms to their utility functions, the Sugarscape simulation can recreate many spatial features and functions of a society and economy, including the emergence of ethnicities, borders, border trade, a segregation of people by their aversion toward pollution, etc. (panel B). It is also possible to extend the simulation with non-local relationships between agents, to make Sugarscape the basis of a social network simulation.
Figure 5: Example output from network models of agents trying to optimize their network positions, adapted from Refs. [13] and [90]. Panel A illustrates an economics-style model where agents optimize a utility function that represents a trade-off between the number of agents whose information can reach them (directly or indirectly through paths of directed links) and the in-degree (i.e., the number of direct-link neighbors) [13]. The figure shows agents updating their links until a Nash equilibrium (which is an absorbing state) is reached (at time step 11). The updating proceeds by agents either configuring their network to what would have been optimal in the previous time step (green arrows) or, with some probability, doing nothing (blue arrows). Panel B presents snapshots of a model with agents again attempting to optimize conflicting objectives: namely, simultaneously maximizing their closeness centrality (a proxy for power) and minimizing their degree (the cost of communication) [89]. The agents can reconfigure their position locally (within their second neighborhoods) and use reinforcement learning based on their own and their neighbors’ performance to decide their updating strategy. The updating strategies are codified as a priority list of actions to take (e.g., deleting the link to the lowest-degree neighbor). The symbols represent actions of highest priority for the respective agent.
6 Complexity science and game theory
Around the mid-1980s, another research field, complexity science, appeared on the scene. Although it thematically overlaps with systems science, cybernetics, etc., complexity science had, at the outset, other sources of inspiration: chaos theory, fractals, cellular automata, and a shared sense of the importance of interdisciplinary research [212, 88]. There were (and still are) deeper differences between systems and complexity science. Systems science, on the one hand, primarily focuses on closed systems with irreducibly complicated interactions (e.g., autopoiesis in organizational engineering [179]). Complexity science, on the other hand, stresses complex patterns emerging [87] from simple rules (like, say, Conway’s “game of life” being capable of universal computation [168]). The way emergence has been conceptualized in complexity science since the mid-1990s has deep connections to the social sciences [80], being akin to Thomas Schelling’s and James Coleman’s micro-to-macro transitions [35, 181, 153], and F. A. Hayek’s “spontaneous order” [75].
6.1 Agent-based models
While complexity science has always had one foot (out of more than two) in the social sciences [^5] For example, Brian Arthur’s complexity economics [7] or urban scaling theory [20]., its relevance for our story mostly comes from agent-based models (ABMs). These were not invented in complexity science—for example, some of the social simulation models mentioned above are ostensibly agent-based—but complexity science was the first discipline to treat agent-based modeling as a comprehensive framework connecting micro- and macro-level phenomena.
Traditionally, the behavioral responses of agents have been encoded by simple decision rules [63, 62]. Although always considered unorthodox, ABMs have found their way into relatively conventional social science [117, 177]. A notable example is their use in analytical sociology [77, 78] and Epstein and Axtell’s more interdisciplinary Sugarscape simulations [44] (see Fig. 4). Economics has also seen a use of AMBs [46, 110], perhaps most frequently in the computational finance literature [108]. These studies include the use of reinforcement learning neural-network agents [18]. Finally, management science, which was keen to import complexity science concepts such as the “edge of chaos” [157], ventured into the realm of ABMs as well [147], especially to study innovation [111, 53] with the so-called NK model of evolution [99].
6.2 Game-theoretic agents
The use of ABMs in political science [9], economics [10], and, to some extent, anthropology [11] usually rests on a game-theoretical foundation [121] assuming people make decisions by maximizing utility functions. When different time horizons clash in the agents’ utility-optimization, we arrive at a social dilemma—a situation where the agents are faced with a choice to behave selfishly (typically leading to higher profits in a short perspective) or prosocially (engaging in cooperative efforts to build society). The study of human cooperation from this perspective has evolved in peculiar ways, where different sub-communities of the field have split and later merged again. In the 1980s, the inspiration came from population biology [123], with the sedentary transition of human prehistory as the explanandum [25]. In the early 1990s, the mathematics of complexity emerging from simplicity was a main driving force of the research [148]. In parallel to these developments, social dilemmas have been studied in economics in several waves. Notably, starting from Fehr et al.’s works on punishment and public goods [47], the early 2000s saw an increased interest in behavioral economics experiments about free riding and the over-exploitation of shared resources.
6.3 Artificial life
Yet another, wildly different field that spawned from complexity science and is of relevance to us is artificial life [112]. This field—with roots in John von Neumann’s search for self-replicating cellular automata [142]—attempts to create computer programs with properties analogous to those of living organisms. Although artificial life rarely deals with higher-order cognitive processes, the initial hype around a series of conferences organized by Chris Langton at the Santa Fe Institute [105] was highly reminiscent of today’s debates about the hypothesized consciousness of LLMs and the purported impending singularity induced by “artificial general intelligence” [45]:
Within fifty to a hundred years, a new class of organisms is likely to emerge. These organisms will be artificial in the sense that they will originally be designed by humans. However, they will reproduce, and will evolve into something other than their initial form; they will be “alive” under any reasonable definition of the word. These organisms will evolve in a fundamentally different manner than contemporary biological organisms, since their reproduction will be under at least partial conscious control, giving it a Lamarckian component. The pace of evolutionary change consequently will be extremely rapid. The advent of artificial life will be the most significant historical event since the emergence of human beings. The impact on humanity and the biosphere could be enormous, larger than the industrial revolution, nuclear weapons, or environmental pollution. We must take steps now to shape the emergence of artificial organisms; they have potential to be either the ugliest terrestrial disaster, or the most beautiful creation of humanity.
Thirty years later, the field of artificial life still studies similar questions, still primarily in silico, but is somewhat more modest in its future predictions [58].
6.4 Network science and networking agents
Representing social relations between individuals or institutions as graphs has a long, albeit patchy, history in the social sciences [52]. The basic tenets were outlined in the 1930s by Helen Jennings and Jacob Moreno, and embrace de Saussurean structuralism in its recursiveness: actors’ roles are determined by how they are connected to other actors [140]. This means that if one maps out the network of connections, there should be mathematical ways of characterizing the nodes’ roles by measuring their positions in the network. Much of social network analysis [216] and network science [14] builds on this idea and thus seeks to develop measures linking positions and roles, structure and function.
The works of Milgram [205], Granovetter [64], and others in the 1960s and 70s led to the insight that people are remarkably good at using their social networks [29]. Along with the above-mentioned assumption that network positions determine people’s opportunities, and the economic dogma that people strive to optimize their prospects, it follows that economic agents should seek to improve their network positions [178, 13, 89]. Figure 5 shows two models of agents trying to optimize their network positions using different utility functions and updating algorithms. The model in Fig. 5A adopts an economics-style approach, assuming that agents can calculate their optimal actions. In contrast, the agents of Fig. 5B learn their networking strategies through reinforcement learning. Note that the network structure of both these models fluctuates quite wildly compared to social networks in the wild, suggesting a model-reality gap to close in future studies.
7 The age of big data
Around the first years of the new millennium, the AI winter gave way to a spring and summer that still lingers. The one-sentence summary of this development is: By using evermore resources, general methods can vastly outperform earlier, specialized ones [71]. The period up to today can be divided by the advent of near-human-level chatbots, where this section covers the first (and epistemically more disruptive) part. During this period, there was a concomitant advancement in hardware, data collection, and analysis methods, accompanied by growth in the information economy and an increased acceptance of the deployment of algorithmic decision making in areas such as insurance and corporate hiring [151]. This was fed into a feedback loop that catalyzed societal changes as interactions increasingly occurred online, thereby producing more data to fuel the algorithms. The zeitgeist is well captured by the McKinsey Global Institute’s 2011 report touting big data as “the next frontier for innovation, competition, and productivity” [119].
The idea of extracting insights from vast streams of data—articulated decades earlier by the likes of Stafford Beer (Fig. 2) and Marshall McLuhan (“information overload = pattern recognition” [127])—was finally becoming reality [196]. Around 2010, data science and data scientist became the established terms for this new guild and its members.
7.1 Advancements in (un)supervised learning
The big-data era did not happen so much because of breakthroughs in machine learning (ML) algorithms as the ability to scale old ones to orders-of-magnitude larger datasets. The resurgence and mainstream adoption of deep learning [82, 183, 109] is probably the best example of this. In 2012, AlexNet, a deep-learning architecture developed by Krizhevsky, Sutskever, and Hinton, won an esteemed computer vision challenge by a legendary margin [102]. AlexNet demonstrated the efficacy of graphics processing units for training deep neural networks, which put a renewed emphasis on hardware. Following AlexNet, subsequent key innovations included the development of convolutional neural network architectures such as VGGNet [194, 183], GoogLeNet with inception modules [198], and ResNet, which introduced the residual learning framework that enabled unprecedented network depths [76]. These developments established deep-learning architectures in computer vision with extensive industrial applications in autonomous vehicles, facial recognition, mass surveillence [95], and image-based diagnostics.
Parallel advancements occurred in natural language processing (NLP), marked notably by the introduction of recurrent neural network architectures like “long short-term memory” networks for sequence modeling tasks [84, 183]. This progress invigorated machine translation, speech recognition, and language modeling [65]. The release of Word2Vec by Mikolov et al. [135], which efficiently generated word embeddings that capture semantic relationships, transformed NLP methodologies and laid the groundwork for later embedding-based models. By 2017, the introduction of transformer models, particularly the seminal “Attention is all you need” paper by Vaswani et al. [210], took AI closer to being generative in a meaningful way, paving the way for large language models like GPT and BERT.
7.2 The revival of reinforcement learning
Another ML paradigm that got a revival in the early 2010s was reinforcement learning [197]. This concerns, very briefly, how to iteratively learn optimal decision-making. In particular, the highly publicized Go match between the reinforcement-trained deep-learning architecture AlphaGo and Korean Go master Lee Sedol [189] in 2016 revitalized this by then already mature topic.
Since humans and animals also learn decision-making successively by their mistakes, there are deep connections between reinforcement learning and other areas of science. For example, Thorndike noticed in 1911 that cats find a balance between exploitation and exploration in problem-solving experiments [201]. This dilemma, whether to stick to the best known policy (local optimum) or search for even better ones (global optima), is a recurrent theme in fields as diverse as robotics, optimization, economics [5], management [120], and finance [32].
7.3 Computational social science and digital traces
The initially corporate term “data science” was soon adopted by academia, with universities launching degree programs, associations springing up, journals being founded, etc. Some authors see the field of computational social science as nacent of this development [217, 107, 175], whereas others point to deeper roots [91, 213]. Regardless, toward the turn of the millennium, an increasing availablility of data more or less directly recording human behavior—digital trace data—made the Internet a valuable data source for the social and behavior sciences [66].
Even if there were precursors over a decade earlier [37], research with digital traces became popular around the final years of the millennium. Digital trace datasets are special in the sense that they are not gathered primarily for the sake of science. Moreover, they are often larger than the typical social-science data sets in earlier literature and thus require other analysis methods and different research questions. For these reasons, it is perhaps not so surprising that early papers on digital trace data were typically carried out by people with a background in the natural or formal sciences. As an example of such an early study, Huberman et al. [92] found a power-law in histograms of arrival rates of unique visitors to websites. They went on to explain this observation by a simple random-walk style model, but—as several other early computational social science articles—did not relate it to the bigger social-science context. [^6] E.g., Arthur’s economics of increasing returns [6] was probably applicable to the online platforms of that era.
As more social scientists saw the value of digital traces, the early years of soul searching gave way to a more broadly informed computational social science. By the mid-aughts, the field had matured enough for studies to become more integrated with mainstream social science. Sagalnik, Dodds, and Watts’ study of success and failure at a website for music sharing is a prime example of this development [176].
Figure 6: The DARPA Red Balloon challenge was a contest in social mobilization for geo-location to commemorate the Internet’s 40th anniversary. On October 29, 2009, ten red balloons were positioned at undisclosed locations across the continental USA (panel A). The first participant to correctly report the location of all was awarded \$40,000. The problem was designed to exploit (then) new technologies like social media and crowdsourcing to scale up the recruitment and organization of human work. Panel B illustrates the winning team’s strategy for incentivizing team-member recruitment inspired by Ref. [42]. Discovering a balloon earned a \$2000 share of the total prize sum, while recruiting someone earned half of that person’s award [199].
7.4 Crowdsourcing
Another data source (in addition to digital traces) that widespread online services made possible was crowdsourcing—the mobilization and coordination of people through digital platforms. Crowdsourcing has roots in optimization by competition, like Axelrod’s computer tournaments (to find the optimal strategy for the iterated Prisoner’s Dilemma [9]) and later data mining and machine learning contests like the KDD Cup and the Netflix Prize. It has been nicknamed “artificial artificial intelligence” for taking the distributed, computer-coordinated architecture from AI, but with humans rather than machines as computing nodes.
Another notable type of crowdsourcing challenge involved the planning and implementation of social mobilization. The most prominent example in this category was DARPA’s Red Balloon challenge, where teams attempted to locate ten balloons across the continental US. The winning team used a hierarchical system to distribute the award between people actually reporting the locations of the balloons and those recruiting the reporters [199, 161] (see Fig. 6).
An even more influential form of crowdsourcing-based science was the use of Amazon Mechanical Turk (MTurk) [26, 27]. Designed to recruit and pay workers for small computer-reportable tasks, originally motivated to build training data for image classification, it found use in the social and behavioral sciences as a platform for experiments. Even though MTurk experiments have always been criticized [100]—mostly because it is hard to know who the workers really are—they became very popular, with over a thousand papers per year by the mid-2010s [27].
7.5 Science of prediction
The era of big data brought the insight that AI could predict (at least some kinds of) human behavior extremely well, better even than the theories of domain experts. Even if much could be said about the role of prediction in the scientific endeavor [203], many saw this as a call to reform the scientific method. This sentiment was accentuated by the replication crisis in the behavioral sciences [152], so by the mid-2010s, change was very much in the air. The more extreme voices foresaw an “end of theory” [3], where computer predictions would replace theory building (in all senses of the word). More moderate methodologists, however, took the rising AI prowess as an impetus to reform the methodology of the social and behavioral sciences, with computational social science leading the way [66, 85].
Fast forward to the present and this debate is still reverberating [86, 85], and, yes, there has been a revolution—especially in econometrics, but also in pockets of public health, political science, psychology, and sociology—but by the widespread adoption of causal inference methods [81], rather than validation by prediction. While this development surely has raised the awareness and quality of the scientific explanations coming out of these fields, some of the problems pointed out a decade earlier [215] remain but have fallen out of focus.
Finally, note that prioritizing prediction over understanding (or theorizing) is not limited to applications of ML. For example, the big-data era furthermore saw a shift towards prediction as a criterion for success in discrete choice modeling (in fields like econometrics and marketing) [204, 173].
8 The age of generative AI
When the release of ChatGPT in November 2022 brought an AI chatbot capable of assisting people in work and their everyday lives [93, 21], much of the futurism of the preceding age of big data gave way to something more familiar. A chatbot itself was nothing new: the first successful one was Eliza from 1966, and two years later Parry was claimed to have passed the Turing test [187, 219]. However, ChatGPT didn’t only sound like a human; it said useful things for a wide variety of situations. It thus made the AI’s capabilities accessible, and ostensibly less like interacting with a black box [106] of superhuman predictive powers (although it was). This opened an entire orchard of low-hanging fruit as social and behavioral scientists could simply replace humans with chatbots in their favorite experiments (or surveys, or crowdsourcing tasks), and triggered a frenzy that, as of now, shows no sign of dissipating. Furthermore, tasks such as text analysis and image classification became much more accessible. It is not the intention of this paper to review this literature, but we will sketch some of the trends in this still young era.
8.1 Technological developments
From 2017 onward, the AI research and industry witnessed an exponential growth driven predominantly by transformer-based models and generative AI. The Generative Pre-trained Transformer (GPT) series of models showed substantial improvements in natural language understanding and generation [164]. The Bidirectional Encoder Representations from Transformers (BERT) enabled contextual embeddings and thereby greatly advanced NLP tasks such as question answering and sentiment analysis [41]. These models are the basis for the LLMs that power the text generation in today’s AI agents.
Another milestone was Generative Adversarial Networks (GANs) that demonstrated hitherto unseen levels of realism in image synthesis, style transfer, and data augmentation [98]. Similar success was achieved by so-called diffusion models [83], which have proven to be more stable and are in more common use today for text-to-X (images, video clips, sound, etc.). At the time of writing, LLMs and diffusion models are combined in commercial applications to make the AI multimodal.
Generative AI holds the distinction of being the first type of successful ML platform to violate the “garbage in, garbage out” dictum since they are capable of generating non-factual output even though trained on validated data. This is because their memories are reconstructed on the fly from distributed bits of information rather than retrieved from a fixed storage [172]. Indeed, the earliest LLMs made an impact with their ability to generate coherent but wholly imaginary stories [165]. An increased factuality has later been imposed on them by fine-tuning on curated training data [41, 2].
Market analysts currently believe in generative AI becoming (if not already being) profitable [33]. On the other hand, ever since the launch of ChatGPT, some commentators have seen omens of an approaching plateau in the AI’s capabilities, despite the enormous investments [172, 188]. Several commentators also raise concerns about their resource demand, their unfair use of training data, depriving people of self-expression, non-accountability, etc.
Figure 7: An illustration of the LLM-chatbot-driven simulations of Ref. [155] where 25 agents simulated the inhabitants of a village. Each agent was seeded by a description of their background and relation to other agents. The authors observed several aspects of emergent sociality, such as the coordinated organization of a Valentine’s Day party. (Reprinted with permission from the authors.)
8.2 Active topics
We are approaching the present and thus the end of the scope of this paper. For the sake of the reader, we will sketch the situation as of writing.
Replacing study objects by chatbots
The first, and to some extent, least creative type of sociobehavioral study employing AI chatbots is just to let them replace humans in experiments or surveys. These types of studies primarily have a value for documenting the capabilities of AI at the time of the study [186]. They could also serve as a basis for future benchmarking or potentially replace pilot studies in experimental designs [195, 12, 229]. The face value of these studies may seem minor. After all, the actual study objects are not human. But then again, we are in the chatbot era precisely because AI is good at emulating humans, and this research will contribute to better future models of humans.
In an early paper of this kind, Park et al. demonstrated that communicating LLM agents could organize an event in collaboration and show other signs of emergent sociality [155] (see Fig. 7). Other papers replicated well-established human behavioral studies with AI: for example, Refs. [131, 50] studied a number of game theoretic experiments with LLMs to conclude that machines are slightly more cooperative than humans. Other studies, on the other hand, found them less cooperative [72]. This reflects the general sentiment at the time of writing: AI can emulate humans well enough to replace them in some intellectual activities, but is easily distinguishable from humans in certain tasks [207, 2, 200].
Studying human-AI interactions
If we are to believe the AI industry, we don’t have any alternative but a future with a much larger presence of AI. Understanding the emergent social consequences of this development is an important research direction [206, 158, 104, 184], and the motivation of several studies with AI agents.
Much of the research in this direction so far has focused on the ability of AI to influence people’s opinions. For example, Costello, Pennycook, and Rand [36] demonstrate that dialogues with AI can cause people to relinquish their beliefs in conspiracy theories for a considerable period. In another study, Hackenburg and Margetts [69] found that LLMs can effectively persuade people on political issues. With a careful study design, one could imagine studying AI-rich futures in the wild, but so far, we are only aware of ethically compromised attempts at this [150] (cf. Fig. 1D).
Another type of research on this theme concerns how human behavior changes when interacting or collaborating with machines compared to other humans. For example, Ref. [118] found that human-machine populations need both peer rewards and peer punishments to overcome human anti-machine bias in economic games.
Improving analysis tools
Foundation models like LLMs are capable of many forms of text and image analyses; maybe not as reliably or with the same interpretability as specialized NLP or image analysis tools, but very cheaply and accessibly.
For example, Bermejo et al. [19] note that for some classification tasks, LLMs can be more accurate than human annotators. Krugmann and Hartmann [103] are equally optimistic about LLMs’ capabilities for sentiment analysis [56], where Törnberg [202] gives a primer to social scientists for using LLM chatbots to analyze texts. Furthermore, Karjus [97] lists topic classification, historical event cause analysis, social network inference from literary texts, and missing data augmentation as some tasks AI is capable of aiding social scientists with.
Replacing scientists by chatbots
Replacing people by chatbots is an idea that has not stopped with experiments and media analysis—several studies have sketched how chatbots can replace scientists either for some specific tasks [223, 228] or entirely [227, 114, 28, 214]. Among the partial tasks where AI could potentially replace social and behavioral scientists are conversational interviews [223] and moderation in focus groups [228]. The papers proposing AI scientists covering the entire scientific process are usually focused on natural science or engineering, but are also sufficiently general to seem relevant to the social and behavioral sciences. Some of these papers are more programmatic in nature [227], whereas others have implemented pipelines and proven that automated paper production can pass the scrutiny of human peer review [114, 28].
8.3 The missing topic: Using AI to understand humans
Despite the surge of research with AI agents, and a flood of research-free perspective and opinion papers pointing out the possibilities [67, 195, 12], studies focusing on humans per se are bewilderingly rare. The majority of AI studies coming out today can, at best, be seen as pilot studies for future investigations of actual social behavior. Even agentic-AI papers whose stated motivation is to understand humans—like Argyle et al. [4], who use experiments with AI agents to validate theories about political persuasions—make the relative performance of various AI setups a take-home message. This is especially surprising given that interpreting human behavior, emotions, etc., to provide machines with an actionable input is a mature topic within the field of human-computer interaction [154]. Perhaps this fruit is just not as low-hanging as the other ones listed above and is simply biding its time to be picked.
8.4 Drawbacks
GenAI-based approaches come with some drawbacks that previous methods did not have. For example, both the rapid development and the inherent stochasticity make LLM studies difficult to replicate. This situation is exacerbated by the lack of a unified methodology  [211]. Moreover, the guardrails and fine-tuning put in place to prevent LLMs from acting offensively or leaking private information can stop working when the models are customized [163]. Finally, generative AI is now in the process of overwriting the human textual corpora that used to serve as its training data. This risks deteriorating both future AI [188] and human knowledge [186].
Figure 8: A timeline of some key moments covered in this paper.
9 Conclusions
The 75 years leading up to today’s agentic AI (Fig. 8) tell a story of humans and machines being deeply entwined. Maybe it all goes back to us trying to understand ourselves, for what one cannot create, one does not understand (to paraphrase Richard Feynman). From the very first social simulation studies, via the early AI’s objective of making machines that learn like children [219, 145], to today’s LLM-based behavioral experiments, the criterion for success is that machine behavior has been indistinguishable from that of humans.
Maybe it is the deeply human fascination with replicas of ourselves and our environment [16, 57] that makes us readily accept simulations as a scientific explanation. Indeed, the methodology of simulation seems to have been in place already before scientists had access to the hardware: recall the 1949 analog simulator of Keynesian economics, the Moniac [192], or Helen Abbey’s 1951 stochastic simulations by hand of the Reed-Frost model of epidemics [1]. Indeed, the highest standard of verification has always been the ability to recreate and forecast what we study [85, 86]. This, on the other hand, has not sufficed for computational approaches to become mainstream, like the statistics revolution in the behavioral sciences [230]. It is tempting to speculate that this is because simulations attempt to explain larger chunks of reality than the typical few causal relations representing the discovery of a sociology or psychology article, or the stylized mathematical models of an economics theory.
Until chess programs reached superhuman levels in the 1990s [145], the question of whether machine intelligence should be as good as possible or as human-like as possible made little sense. However, now branches defined by these two goals pull the AI-augmented social and behavioral sciences in different directions. The as-good-as-possible mindset was behind the big data and data-science boom of the 2010s. With new data sets of unprecedented size collected from the digital breadcrumbs we leave behind (thus realizing an old cybernetic dream—cf. Fig. 2), ML was able to predict human behavior in ways unattainable to social and behavioral science. Turning this situation into knowledge creation became the raison d’être of the new field of computational social science, and spurred an epistemic discourse around the role of prediction in science that is still unfolding [85, 86, 218]. Lately, however, obtaining digital trace data for academic research has become increasingly restricted, ushering in the changes described below.
With the advent of ChatGPT in 2022 and other easily accessible LLM-based chatbots, the winds changed. The predominant mindset became to harness AI’s ability to mimick humans—i.e., to strive for AI to be as aligned with human behavior as possible [22, 2, 70]. Naively, this does not seem to require a substantial rethinking of the traditional scientific process. One can simply replace humans with chatbots and proceed with mainstream methods. The catch, of course, is that even though one can align AI and human behavior in some aspects, it is hard to estimate how much this alignment generalizes [2, 200]. Moreover, when it comes to indirect, iterative tasks such as reinforcement learning [113], or planning [96], LLMs seem to be less successful in replicating humans. So, in conclusion, generative AI provides a standardized black box model of humans  [106] that may be practical and convenient but involves many epistemic subtleties [134].
The current fervor of GenAI-fueled social science may subside, but its impact will last. The recent technological progress provides more than just models of humans. We (as in humans and AI combined) now have new ways of processing, summarizing, and structuring data, as well as new methods for generating, systematizing, and disseminating knowledge. Together with the new methods and epistemic perspectives of the big-data era, we now have an abundance of opportunities to understand the social world and human behavior. This will be ever more important since the very same technologies are changing our society and behavior, and this change also contributes to our scientific introspection: “Every new technology allows us to learn something new about ourselves” [15]. Social science has always been special in that it is related to what it studies—a “subject-subject relation” [60]—only now the instruments of inquiry are also part of the equation.